i m unable to solve 1 paragraph type query
a rxn is said to be 1st order if its rate is proportional to the concn of reactant see 1 rxn A--- B+ C at t=0, A=a , b=o, c=0 and at time t A= a-x, , b=x, c=x now see rate of rxn given by dx/dt= k(a-x) and integrated rate eqn for a given rxn is k= 1/t ln(a/a-x) a= initial conc and a-x= concn of a after time t
1- consider a rxn a- ----3b+2c rate cnst= 1.386 * (10)-2 starting with 2 moles of a in 12.5 vessel initially rxn is allowed to take place at cnst pressure and at 298 k what is the concn of b after 100 min
2- thermal decomposition of cpd x is a 1st order rxn 75%of x is decomposed in 100 min how long it will take for 90% of cpd to decompose

Grade:12

1 Answers

Sunil Kumar FP
9 years ago
given reaction is a------3b +2c
k=1.386*10^-2
k=1/t*2.303log(2/2-x)
1.386*10^-2=1/100log(2/2-x)
x=1.49 mole
conc=1.49/12.5=.119mol/lt

(2)
T75%=klog100/25
T90%=klog100/10
T90%=166min